# Thermalization in Krylov Basis

### v1

@@ II.2 Krylov space and nature of thermalization
To proceed, let us first compute Lanczos coefficients for three distinct initial states we have considered in the previous section [^5] Lanczos coefficients for the model under consideration have also been computed in Noh_2021 ; Scialchi:2023bmw ; Espanol:2022cqr ; Trigueros:2021rwj .. The results for $N=10$ are shown in figure 4
@@
Figure 4: Lanczos coefficients $a_{n},b_{n}$ of three initial states $|Y+\rangle,|Z+\rangle,|X+\rangle$. $b_{n}$ and $a_{n}$ are shown with blue and brown circles, respectively. The numerical results are presented for $N=10$ in which the dimension of Krylov space for a generic initial state is about $529$.
@@
Although one could recognize some differences between these three plots, the differences are not significant and essentially the Lanczos coefficients for all three cases have qualitatively the same behavior. Therefore, one may want to conclude that the naive behavior of Lanczos coefficients could not provide a quantity to make a distraction between these three cases.
@@
As a final observation, one notices that at the symmetric axis of $\phi=\pi$, while the effective inverse temperature or the absolute value of the density of energy decreases almost monotonically from $\theta=\frac{\pi}{6}$ to $\theta=\pi$, the variance of $a_{n}$ exhibits a pick around $\theta\approx\frac{5\pi}{6}$ suggesting that the state $|\frac{5\pi}{6},\pi\rangle$ would be among those with the weakest thermalization.
Actually, since we have an exact analytic expression for the energy one can use it as a gauge to validate our numerical results. In fact, by making use of the analytic expression of energy we can compute the density of energy for $\phi=\pi$ slice for arbitrary $N$. In figure 6 we have drown the density of energy for $\phi=\pi$ slice for $N=10$ and 100. We have also presented the first derivative of the density of energy to better explore its behavior.
@@
From this figure, one finds that the density of energy, indeed, exhibits a non-trivial behavior around $\theta=\frac{5\pi}{6}$, though it is not as pronounced as that in the variance of $a_{n}$. Actually, the first derivative is closed but not exactly zero showing that there is no real minimum around the position of the second pick, even though for higher $N$ the situation gets better. Therefore, although the general behavior of the variance of Lanczos coefficients are consistent with the results of other probes, we have seen slightly mismatch at $\phi=\pi$ slice.
Working with Krylov basis, yet another interesting quantity one may consider as a possible measure to probe the nature of thermalization is Krylov complexity (22). Indeed, by making use of the fact that the Krylov complexity is essentially the expectation value of the number operator, one can compute the infinite time average of Krylov complexity Rabinovici:2021qqt and consider it as a possible probe. More precisely, from the expression of Krylov complexity (22) one has
@@
It is then straightforward to compute the infinite time average of complexity for states associated with initial states (6). The numerical result for $N=9$ is depicted in figure 7.
@@
Interestingly enough, the resultant pattern is exactly compatible with those of the absolute value of the density of energy and the effective inverse temperature. In particular, we have found that for states exhibiting strong thermalization the saturation value for complexity is larger.
@@
To get a better understanding of what exactly happens in different points in $\theta-\phi$ plane (initial states), it is useful to explicitly compute the expectation value of local operators to see how the thermalization occurs for different initial states. To do so, we will consider the magnetization in the $z$ direction and compute the following quantity 666 We have also computed the expectation value for the magnetization in the $x$ direction, $S_{x}$, in which we have found that the conclusion is the same as that of $S_{z}$ that is explicitly presented in what follows.
@@
for different values of $\theta$ and $\phi$. The results are depicted in figure 8. Actually, we have computed the corresponding expectation value for 441 initial states [^7] Since the pattern in the figure 5 is symmetric under $\phi\rightarrow 2\pi-\phi$, we have only considered initial states located in $0\leq\phi\leq\pi$. and only few of them have been shown in this figure which are for particular slices given by $\phi=0,\pi$ and $\theta=\frac{\pi}{2},\frac{4\pi}{5},\frac{19\pi}{20}$. The results are relatively compatible with what suggested by the variance of Lanczos coefficients and in exact agreement with what suggested by the infinite time average of complexity.
@@
From the behavior of the expectation value of $S_{z}$ one observes that states with weak thermalization are mostly located around $\theta\approx 0$ while those of strong thermalization are around $\theta\approx\pi$, besides those located near ring of zero $\beta$. Thus, as one moves from $\theta=0$ towards $\theta=\pi$, generally, states show stronger thermalization. Interestingly enough, one finds that the initial states $|Y+\rangle,|Z+\rangle$ and $|X+\rangle$ are not among those states with the strongest or weakest thermalization. It is in contrast with the general belief that states with strong thermalization are located near the ring of zero $\beta$. In particular, states $|Z+\rangle=|0,\phi\rangle$ and $|Z-\rangle=|\pi,\phi\rangle$, for arbitrary $\phi$, exhibit weak and strong thermalization, respectively.
@@
To be more concrete let us look at $\phi=0$ slice where we have presented results for $\theta=0,\frac{\pi}{5},\frac{\pi}{2},\frac{3\pi}{5},\frac{4\pi}{5},\frac{19\pi}{20},\pi$ in figure 8. By making use of these results one finds that at $\theta=0$ we have weak thermalization though it becomes stronger as one approaches the ring of zero $\beta$. Then it becomes weaker as one moves further forwards $\frac{\pi}{2}$ after which it keeps becoming stronger all the way to $\theta=\pi$. Interestingly enough, we have found that the behavior is consistent with the behavior of the infinite time average of complexity which is also in qualitative agreement with the behavior of the variance of Lanczos coefficients.
For $\theta=\frac{\pi}{2}$ slice, one finds very weak thermalization at $\phi=0$ and then it becomes very strong at $\theta=\frac{\pi}{2}$ where the slice intersects the ring of zero $\beta$. Then, it becomes weaker as one moves towards $\theta=\pi$ point. For $\theta=\frac{4\pi}{5}$ slice, we get strong thermalization almost on the whole slice, though it shows slightly stronger behavior in a point where the slice intersects the region of zero $\beta$. Of course, they become relatively weaker as one moves towards $\phi=\pi$. Finally, for $\theta=\frac{19\pi}{20}$ slice one has always strong thermalization all the way from $\phi=0$ to $\phi=\pi$.
An interesting observation we have made is that, unlike the $\theta=\frac{\pi}{2},\frac{4\pi}{5}$ slices where we see a significant change at the ring of zero $\beta$, for $\theta=\frac{19\pi}{20}$ slice the change is not so significant, in agreement with the behavior of the variance and the infinite time average of complexity.
@@
Looking at $\phi=\pi$ slice, from the expectation value of $S_{z}$ we find that the weakest thermalization occurs at $\theta\approx\frac{\pi}{6}$ while it becomes relatively stronger as we move towards $\theta=\pi$. Surprisingly, we did not observe any further special point in this slice in agreement with the the behavior of the infinite time average of complexity and in contrast to the behavior suggested by the variance in which we would expect to have a state with relatively weaker thermalization around $\theta\approx\frac{5\pi}{6}$.
II.3 Inverse participation ratio
It is worth mentioning that the nature of weak and strong thermalization for certain XY Ising model has been studied inPrazeres:2023hce by making use of the inverse participation ratio Short:2011pvc . Thus it is worth looking at this quantity for our model too.
Consider a state whose expansion in the energy eigenstates is $|\psi\rangle=\sum_{i=1}^{\cal D}c_{i}|E_{i}\rangle$, where $c_{i}=\langle E_{i}|\psi\rangle$. Then, the inverse participation ratio is defined by
$\lambda=\frac{1}{\sum_{i=1}^{\cal D}|c_{i}|^{4}}\,,$ (26)
that is essentially a quantity that measures the number of energy eigenstates that contribute to the state $|\psi\rangle$. Note that $1\leq\lambda\leq{\cal D}$. In fact, when only one energy eigenstate contributes to the state the inverse participation number is one, while when all energy levels equally contribute to the state it is equal to ${\cal D}$.
Now let us compute the inverse participation ratio for general initial state (6). The result is depicted in figure 9. Note that what is drown in this figure is the logarithm of the inverse participation ratio, $\ln\lambda$. Interestingly enough, one observes a significant correlation with all quantities we have considered so far, including the variance of Lanczos coefficients.
It is worth noting that the behavior of the inverse participation ratio is in perfect agreement with that of the infinite time average of complexity. It shows that the saturation value for complexity is greater for states consisting of more energy eigenstates Scialchi:2023bmw ; Rabinovici:2021qqt .
It is important to note that using the behavior of the effective inverse temperature and normalized energy it was suggested that the weak thermalization occurs for the states for which the expectation value of the energy is near the edge of the energy spectrum Banuls:2011vuw ; Sun:2020ybj ; Chen:2021 ; Lin:2016egw . Here, we have seen that whether a state exhibits weak or strong thermalization is correlated to its inverse participation ratio as suggested in Prazeres:2023hce . More precisely, the nature of the thermalization of a state is closely related to the number of energy eigenstates that contribute to the state. A state consisting of more energy eigenstates likely exhibits stronger thermalization. Note that our computations of the expectation value of local operators confirm this behavior too.
It is also interesting to see that looking at $\phi=\pi$ slice, we get similar behavior as that in the variance of Lanczos coefficients, Namely, there are two picks around $\frac{\pi}{6}$ and $\frac{5\pi}{6}$. Unlike the variance of Lanczos coefficients in this case one can see that although the first pick represents the state with the weakest thermalization, the second pick is the artificial of the finite $N$ effect. To see this, we have presented the logarithm of the inverse participation ratio for $N=10$ and $N=11$ in figure 10 from which it is evident that the second pick is removed as we go to higher $N$. Interestingly enough, the correction due to higher $N$ done not significantly alter other features of the inverse participation ratio.
Figure 10: The logarithm of inverse participation ratio of $\phi=\pi$ slice for $N=10,11$. One observes that the second pick (minimum) is removed as one goes to higher $N$.
It is also interesting that for $N=11$ the behavior of inverse participation ratio at $\phi=\pi$ is very closed to that of the density of energy (see figure 6).
III Discussions
@@
In this paper, we have studied thermalization for a closed quantum system using the Krylov basis. Actually, our main motivation to do so is that, by definition, under time evolution a quantum state propagates over a subspace of Hilbert space known as Krylov space. An advantage (at least theoretically) of working in this space is that we will have to deal with a space whose dimension is usually smaller than the dimension of the Hilbert space.
@@
On the other hand, working within the context of Krylov space, one will have to compute matrix elements of local operators in the Krylov basis. It is then natural to expect that a similar concept may also show up in this context. Indeed, by making use of an explicit example we have shown that the matrix elements of local operators satisfy a similar condition as that in ETH.
We note, however, that unlike in the energy eigenstates matrix elements in Krylov basis are generally complex numbers. Therefore, in this case the situation is more complicated than ETH. Indeed, what we have shown is that the real part of the matrix elements follows a similar behavior as that of ETH and thus we called the statement as Krylov basis thermalization hypothesis (KTH).
For a given operator the contribution of the imaginary part is in the off diagonal terms that may cause an oscillatory behavior over thermal equilibrium value which could affect the natural of the thermalization. It is interesting to understand this point betterwork .
We note that the nature of thermalization depends on two important ingredients: the Hamiltonian of the system and the initial state. As we have seen, by construction, the Krylov basis and Lanczos coefficients contain information about these ingredients making them nature objects to probe the nature of thermalization.
We have shown that the infinite time average of Krylov complexity could provide a measure to probe the nature of thermalization. We have seen that the behavior of the infinite time average of complexity is in perfect agreement with the behaviors of the effective inverse temperature and the density of energy. In particular, an initial state exhibiting strong thermalization has relatively larger value for complexity saturation.
@@
To further explore the thermalization properties of the model under consideration we have also evaluated the inverse participation ratio for general initial states. We have observed that strong thermalization occurs for states with relatively greater inverse participation ratio. In other words, a state consisting of more energy eigenstates is more likely to exhibit stronger thermalization. We have seen that there is a direct relation between the behaviors of the infinite time average of complexity and the inverse participation ratio.
To verify our proposal we have also computed the time dependence of the expectation value of local operators to explicitly probe the nature of thermalization for the generic initial state given by (6). The results, indeed, confirm our observation based on the behaviors of the infinite time average complexity, the inverse participation ratio and the variance of Lanczos coefficients.
@@
One may also explore thermalization and its nature for this model by evaluating different quantities such as the effective inverse temperature and the infinite time average of complexity. Doing so, one can see that the results are consistent with the behavior of the expectation of the energy (28), that also confirms our expectations. We note that the inverse participation ratio for this model has been studied in Prazeres:2023hce .
@@
Defining an inner product in the space of operators acting on the Hilbert space, one can construct Krylov basis for the operator starting with an initial operator ${\cal O}$. The first element is identified with the initial operator $O_{0}={\cal O}$ (which we assume to be normalized with respect to the inner product) and the other elements may be constructed recursively as follows
@@
where ${\cal O}_{i}$ is defined in (7). It is worth noting that since this initial state (6) is the eigenstate of the above operator, in this case we are essentially studying the time evolution of density matrix associated with the initial state ${\cal O}_{\theta,\phi}=\rho(\theta,\phi)=|\theta,\phi\rangle\langle\theta,\phi|$.
One can also study the variance of Lanczos coefficients $\hat{b}_{n}$ associated with the initial density matrix. Doing so, one finds the corresponding variance results in the same conclusion as that for the state studied in the previous section. An interesting observation we have made is that the behavior of variance $\hat{b}_{n}$ in operator growth is actually identical with that obtained from $a_{n}$ in state growth. It would be interesting to understand this point betterwork .

### v3

@@ III.1 Variance of Lanczos coefficients
To explore this idea, let us begin by calculating the Lanczos coefficients for three different initial states that we have discussed in the previous section. [^10] Lanczos coefficients for the model under consideration have also been computed in Noh_2021 ; Trigueros:2021rwj ; Espanol:2022cqr ; Bhattacharya:2023zqt ; Bhattacharya:2022gbz ; Scialchi:2023bmw .. The results for $N=10$ are shown in figure 8
@@
Figure 8: Lanczos coefficients $a_{n},b_{n}$ of three initial states $|Y+\rangle,|Z+\rangle,|X+\rangle$. $b_{n}$ and $a_{n}$ are shown with blue and brown circles, respectively. The numerical results are presented for $N=10$ in which the dimension of Krylov space for a generic initial state is about $528$.
@@
Although one could recognize some differences among these three plots, the differences are not substantial. Essentially, the Lanczos coefficients exhibit qualitatively similar patterns across all three cases. As a result, one might conclude that the straightforward behavior of Lanczos coefficients may not offer a distinct metric to differentiate between these three cases.
@@
Notably, along the symmetric axis at $\phi=\pi$, while the effective inverse temperature and the absolute value of the energy density decrease almost monotonically from $\theta=\frac{\pi}{6}$ to $\theta=\pi$, the variance of $a_{n}$ exhibits a minimum around $\theta\approx\frac{5\pi}{6}$. This suggests that the state $|\frac{5\pi}{6},\pi\rangle$ is among those with the weakest thermalization.
To validate our numerical results, we can leverage the exact analytic expression for the energy. This allows us to compute the energy density for the $\phi=\pi$ slice for arbitrary $N$. In figure 10, we present the energy density for $N=10$ and $N=100$ to further examine its behavior.
@@
From this figure, we observe that the energy density exhibits non-trivial behavior around $\theta=\frac{5\pi}{6}$, though this is less pronounced than the variance of $a_{n}$. Using the explicit form of the energy density for large $N$ given in equation (9), we can compute its derivative with respect to $\theta$ for the $\phi=\pi$ slice
$E^{\prime}=\frac{N}{2}\left(\sin\theta+4\cos\theta\sin\theta-\frac{21}{10}\cos\theta\right)\,,$ (42)
where the prime indicates differentiation with respect to $\theta$. The potential minima can be identified by solving the equation
$\sin\theta+4\cos\theta\sin\theta-\frac{21}{10}\cos\theta=0\,.$ (43)
It is straightforward to see that this equation has only one solution for $\theta\in[0,\pi]$. Therefore, while the second point may appear significant, there is no true minimum around $\theta=\frac{5\pi}{6}$. Remarkably, for large $N$, the position of the minimum remains independent of $N$.
In conclusion, we find a correlation between the behavior of the effective inverse temperature and the variance of the Lanczos coefficients. However, notable discrepancies persist between the two, suggesting that the variance of the Lanczos coefficients may not be a reliable indicator of the nature of thermalization, despite its potential to capture certain aspects of it.
III.2 Infinite time average of Krylov complexity
Working with Krylov basis we note that there is rather a special operator in Krylov space whose matrix elements are proportional to Kronecker delta. More precisely, consider the number operator defined by
${\cal N}=\sum_{n=0}^{{\cal D}_{\psi}-1}\,n|n\rangle\langle n|,$ (44)
that is obviously diagonal in Krylov basis, ${\cal N}_{nm}=n\delta_{nm}$. We note that the expectation value of the number operator, actually, computes Krylov complexity Parker:2018yvk  [^11] See also Avdoshkin:2019trj ; Dymarsky:2019elm ; Dymarsky:2021bjq ; Kim:2021okd ; Adhikari:2022whf ; Camargo:2023eev ; Huh:2023jxt for related works.
that saturates at very late times where the Lanczos coefficients vanish Rabinovici:2022beu . The Krylov complexity is an interesting quantity which relies on both the initial state and the Hamiltonian, akin to Lanczos coefficients.
The evolution of maximally entangled states in the Krylov basis has been studied in Erdmenger:2023wjg , revealing that the growth and subsequent saturation of Krylov complexity are common features of many-body systems, regardless of their chaotic or integrable nature. As a general consequence, we would expect to observe the following behavior at late times
${\cal C}=\overline{{\cal C}}+{\rm small\;fluctuations}\,,$ (46)
where $\overline{{\cal C}}$ is the infinite time average of the Krylov complexity given by
@@ III.2 Infinite time average of Krylov complexity
It was conjectured that one could probe the system’s dynamics by examining the infinite time average of Krylov complexity, with chaotic models exhibiting higher values Rabinovici:2021qqt . In fact, in an explicit Ising model, it has been demonstrated in Rabinovici:2021qqt that the infinite time average of the Krylov complexity increases as one transitions from an integrable model to a chaotic one. However, it is important to note that in addition to the dynamics the infinite time average of Krylov complexity also depends on the initial states. In fact, the infinite time average of Krylov complexity for chaotic systems may or may not exceed that of integrable models Scialchi:2023bmw ; Trigueros:2021rwj ; Espanol:2022cqr .
We have leveraged this insight to propose the infinite time average of Krylov complexity as a measure to probe the nature of thermalization. It is straightforward to compute the infinite time average of complexity for states associated with initial states (6). The numerical result for $N=9$ is depicted in figure 11.
@@
Interestingly, the resulting pattern aligns perfectly with the absolute value of the energy density and the effective inverse temperature. Specifically, we have observed that states exhibiting strong thermalization have a higher complexity saturation value than those with weak thermalization. In conclusion, for the model described in equation (1), we find that as thermalization intensifies, the complexity saturation value also increases. We hypothesize that similar conclusions may apply to generic models.
III.3 Inverse participation ratio
It is worth mentioning that by making use of the inverse participation ratio Short:2011pvc the nature of weak or strong thermalization of certain XY Ising model has been studied inPrazeres:2023hce . Thus it is worth looking at this quantity for our model too.
Consider a state whose expansion in the energy eigenstates is $|\psi\rangle=\sum_{\alpha=1}^{\cal D}c_{\alpha}|E_{\alpha}\rangle$, where $c_{\alpha}=\langle E_{\alpha}|\psi\rangle$. Then, the inverse participation ratio is defined by
$\lambda=\frac{1}{\sum_{\alpha=1}^{\cal D}|c_{\alpha}|^{4}}=\frac{1}{{\rm Tr}(\rho_{DE}^{2})}\,,$ (48)
which is essentially a quantity that measures the number of energy eigenstates that contribute to the state $|\psi\rangle$. Note that $1\leq\lambda\leq{\cal D}$. In fact, when only one energy eigenstate contributes to the state the inverse participation number is one, while when all energy levels equally contribute to the state it is equal to ${\cal D}$.
It is worth also noting that the inverse participation ratio may be given in terms of the infinite time average of the wave function $\phi_{0}$
[…]
Now, let us compute the inverse participation ratio for the general initial state (6). The results are depicted in figure 12, which displays the logarithm of the inverse participation ratio, $\ln\lambda$. Interestingly enough, there is a strong correlation with all the quantities we have examined thus far, including the variance of the Lanczos coefficients.
It is important to highlight that analyses of the effective inverse temperature and normalized energy suggest weak thermalization occurs for states near the edge of the energy spectrum Banuls:2011vuw ; Sun:2020ybj ; Chen:2021 ; Lin:2016egw . In our investigation, we found a correlation between whether a state exhibits weak or strong thermalization and its inverse participation ratio, as indicated in Prazeres:2023hce . Specifically, the nature of a state’s thermalization is closely linked to the number of energy eigenstates contributing to it; states comprised of more energy eigenstates tend to display stronger thermalization. Our computations of the expectation values of local operators further confirm this behavior.
Moreover, the behavior of the (log) inverse participation ratio aligns perfectly with the infinite time average of complexity. This indicates that the saturation value of complexity is higher for states composed of a greater number of energy eigenstates.Scialchi:2023bmw ; Rabinovici:2021qqt .
It is intriguing to observe that when examining the $\phi=\pi$ slice, we see a behavior similar to the variance of the Lanczos coefficients. Specifically, there are two distinct minima located around $\frac{\pi}{6}$ and $\frac{5\pi}{6}$. However, it is important to note that the first minimum corresponds to the state with the weakest thermalization, while the second minimum is an artifact of the finite $N$ effect. To illustrate this, we present the logarithm of the inverse participation ratio for $N=10$ and $N=11$ in figure 13, which clearly shows that the second minimum disappears as we increase $N$. Interestingly, the corrections associated with larger $N$ do not significantly alter the other characteristics of the inverse participation ratio.
Figure 13: The logarithm of inverse participation ratio of $\phi=\pi$ slice for $N=10,11$. One observes that the second peak (minimum) is removed as one goes to higher $N$.
To conclude we note that for $N=11$, the behavior of the inverse participation ratio at $\phi=\pi$ closely resembles that of the energy density, with only one true minimum present; the second minimum is not genuine (see figure 10).
III.4 Time dependent of expectation value
So far, we have investigated several measures to probe the nature of thermalization. It is important to note that our understanding of which states exhibit strong or weak thermalization is derived from energy behavior, or equivalently, from the infinite time average of complexity and the inverse participation ratio, as shown in the figures 1, 11 and 12, respectively.
From these figures, we observe that the state associated with $\theta=0$ (arbitrary $\phi$), corresponding to $|Z+\rangle$, exhibits weak thermalization. In contrast, the state at $\theta=\pi$ (arbitrary $\phi$), corresponding to $|Z-\rangle$, demonstrates strong thermalization. As $\theta$ transitions from 0 to $\pi$, a non-trivial behavior emerges, which is evident in the aforementioned figures.
@@ III.4 Time dependent of expectation value
To validate this expectation, we can compute the expectation value of a typical operator across different states to determine if they exhibit strong or weak thermalization. In fact, to get a better understanding of what actually happens in different points in the $\theta-\phi$ plane (initial states), we will consider the magnetization in the $z$ direction and compute the following quantity 121212 We have also computed the expectation value for the magnetization in the $x$ direction, $S_{x}$, in which we have found that the conclusion is the same as that of $S_{z}$ that is explicitly presented in what follows.
@@
for different values of $\theta$ and $\phi$. In examining the behavior of this expectation value, we note distinct characteristics for strong and weak thermalization. In the strong case, we observe a swift relaxation characterized by a quick rise or fall of the expectation value, followed by a saturation phase around the thermal value, with minor fluctuations. Conversely, weak thermalization shows oscillatory behavior from the outset, oscillating around the thermal value.
To quantify these behaviors, we calculate the ratio of the oscillation size (variance of the oscillation) to the amplitude of the first peak or trough following relaxation. This ratio serves as an indicator of thermalization strength: for states exhibiting weak thermalization, the ratio approaches one, whereas for states with strong thermalization, it is significantly less than one. By assessing this ratio for a specific initial state, we can determine its tendency toward either strong or weak thermalization. For instance, for the state $|y+\rangle$, which is expected to exhibit strong thermalization, the ratio is about 0.01, while for the states $|Z+\rangle$ and $|Z+\rangle$, it is about 0.5 and 0.7, respectively .
By exploring various initial states, we have found perfect agreement with our expectations based on the behavior of energy. Actually, we have computed the corresponding expectation value for 441 initial states [^13] Since the pattern in the figure 9 is symmetric under $\phi\rightarrow 2\pi-\phi$, we have only considered initial states located in $0\leq\phi\leq\pi$. and only few of them have been shown in this figure 14 which are for particular slices given by $\phi=0,\pi$. The results are relatively compatible with what suggested by the variance of Lanczos coefficients and in exact agreement with what is suggested by the infinite time average of complexity.
@@
To further explore this point let us look at $\phi=0$ slice where we have presented results for different values of $\theta$ in figure 14. By making use of these results we find weak thermalization at $\theta=0$, strong thermalization around $\theta=\pi/4$, weak thermalization again at $\theta=\pi/2$, and finally strong thermalization as we approach $\theta=\pi$. Interestingly enough, we have found that the behavior is consistent with the behavior of the infinite time average of complexity.
Looking at $\phi=\pi$ slice, from the expectation value of $S_{z}$ we find that the weakest thermalization occurs at $\theta\approx\frac{\pi}{6}$ while it becomes relatively stronger as we move towards $\theta=\pi$. Actually evaluating the energy expectation value, one can see that the initial state $|\frac{\pi}{6},\pi\rangle$ is very close to an eigenstate of the Hamiltonian. Thus being localized in energy eigenstates one observes an oscillatory behavior for typical operators. This can also be seen from the inverse participation number in which for this state one has $\lambda\approx 1$.
@@
It is worth also noting that we did not observe any further special point in this slice in agreement with the behavior of the infinite time average of complexity and in contrast to the behavior suggested by the variance in which we would expect to have a state with relatively weaker thermalization around $\theta\approx\frac{5\pi}{6}$.
IV Discussions
@@
In this paper, we have studied thermalization for a closed quantum system using the Krylov basis. Actually, our main motivation to do so is that, by definition, under time evolution a quantum state propagates over a subspace of Hilbert space known as Krylov space. An advantage (at least theoretically) of working in this space is that we will have to deal with a space whose dimension is usually smaller than the dimension of the full Hilbert space.
@@
On the other hand, working within the context of Krylov space, one will have to compute matrix elements of local operators in the Krylov basis. It is then natural to expect that a similar concept may also show up in this context. Indeed, by making use of an explicit example we have shown that the matrix elements of local operators satisfy a condition analogous to that of ETH. More precisely, We have demonstrated that for thermalization to occur, the matrix representation of typical local operators in the Krylov basis should exhibit a specific tridiagonal form with all other elements in the matrix being exponentially small.
We have also studied the nature of thermalization in this framework by introducing certain metrics to probe whether a given initial state exhibits weak or strong thermalization. To do so, We have observed that the nature of thermalization depends on two crucial factors: the system’s Hamiltonian and the initial state. The Krylov basis and Lanczos coefficients, by construction, contain information about these elements, making them capable for studying the process of thermalization.
We have shown that the infinite time average of Krylov complexity could provide a measure to probe the nature of thermalization [^14] We also note that a certain state dependence of Krylov complexity has been studied in Kundu:2023hbk which is in perfect agreement with the behavior of the effective inverse temperature and the density of energy. In particular, an initial state exhibiting strong thermalization has relatively larger value for complexity saturation.
@@
To further explore the thermalization properties of the model under consideration we have also evaluated the inverse participation ratio for general initial states. We have observed that strong thermalization occurs for states with relatively greater inverse participation ratio. In other words, a state consisting of more energy eigenstates is more likely to exhibit stronger thermalization. We have seen that there is a correlation between the behaviors of the infinite time average of complexity and the inverse participation ratio.
To verify our proposal we have also computed time dependence of the expectation value of local operators to explicitly probe the nature of thermalization for the generic initial state given by (6). The results, indeed, confirm our observation based on the behaviors of the infinite time average complexity, the inverse participation ratio and the variance of Lanczos coefficients.
@@
One may also explore thermalization and its nature for this model by evaluating different quantities such as effective inverse temperature and infinite time average of complexity. Doing so, one can see that the results are consistent with the behavior of the expectation of the energy (52), that also confirms our expectations. We note that the inverse participation ratio for this model has been studied in Prazeres:2023hce .
@@
Defining an inner product in the space of operators acting on the Hilbert space, one can construct the Krylov basis for the operator starting with an initial operator ${\cal O}$. The first element is identified with the initial operator $O_{0}={\cal O}$ (which we assume to be normalized with respect to the inner product) and the other elements may be constructed recursively as follows
@@
where ${\cal O}_{i}$ is defined in (7). It is worth noting that since this initial state (6) is the eigenstate of the above operator, in this case, we are essentially studying the time evolution of density matrix associated with the initial state ${\cal O}_{\theta,\phi}=\rho(\theta,\phi)=|\theta,\phi\rangle\langle\theta,\phi|$.
One can also study the variance of Lanczos coefficients $\hat{b}_{n}$ associated with the initial density matrix. Doing so, one finds the corresponding variance results in the same conclusion as that for the state studied in the previous section. An interesting observation we have made is that the behavior of variance $\hat{b}_{n}$ in operator growth is actually identical with that obtained from $a_{n}$ in state growth. It would be interesting to understand this point better.